OpenStudy (anonymous):

Can someone help me solve these word problems? The sum of two numbers is 8 less than twice the first number. Their difference is 4 less than twice the second number.

OpenStudy (timaashorty):

As you read this word problem try to pronounce it out. The SUM (+) is 8 > 2(1st number). Lets say the first number is x. The difference (-) is 4 < 2( 2nd number) Lets the say the second number is y. So we'll write th equations like this: a+x=2a - 8 a-x= 2x - 4 Subract 2 a from both sides of the first equation= . x-a= -8 add x-a=-8 by the second equation. + 12 to both sides. 2x=8+4 2x=12 x=6 we end up with 6-a=-8 now we have x-8=-8 & 6-8=-8 add 8 to both of the sides which we'll get a = 14. now just plug in the answers to make sure. a = 14 & x = 6